Shift summary for eve., 10/20/99

Ed Blucher (blucher@hep.uchicago.edu)
Wed, 20 Oct 1999 22:44:42 -0500


Shift Summary for Eve Shift, 10/20/99

On Shift: Ed + Patty,Sydney

Description of shift: Y2K test preparations

We continued work on debugging DAQ in preparation for tomorrow's Y2K test.
The readout is now working, and we've tested two new configurations for
the Y2K test: COSMICS_Y2K, LASER_SCAN_Y2K

Our current plan is to set the clocks forward to January 2000 at 9AM tomorrow
morning. The required tests for the DAQ system should be completed during
the day tomorrow.